OpenStudy (love10129151):

Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a. passes through (-1, 4) and (-5, 2) Equation: Explain why you chose the form you used. b. slope 2, y-intercept -4 Equation:

OpenStudy (love10129151):

Task 2 Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a. passes through (-1, 4) and (-5, 2) Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________ b. slope 2, y-intercept -4 Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________ c. has an x-intercept of 6 and a y-intercept of 3 Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________ d. passes through (1, 2) with slope -5/3 . Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________

OpenStudy (anonymous):

use the formula for finding the slope \(\Large{m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}\) then use the equation of the line \(y-y_{1}=m(x-x_{1})\)

OpenStudy (love10129151):

can you give me a answer

OpenStudy (gorv):

a .......> y-4=(-1/2)(x+1) y=-x/2-1/2 + 4 y=-x/2+7/2

OpenStudy (gorv):

general equation of line y=mx+c m= slope c=y intercept y=2x-4

OpenStudy (gorv):

@love10129151

OpenStudy (gorv):

in a part points are given so we can find slope and line passing through the point are given so we used point slope form